# Activated solids: Spontaneous deformations, non-affine fluctuations, softening, and failure

### v1

@@ III.1 Scaling analysis
Figure 2: (a) Evolution of the global non-affine parameter $X(t)$ at Λ=4, $\Lambda=4,\,10$. (b) Plot of $\langle\Delta X\rangle/a_{\rho}$ versus $\Lambda$ for $N=4096$ particles at varying densities $\rho$, where $\Delta X=X(\Lambda)-X_{eq}$, and $X_{eq}$ is the equilibrium value of global non-affinity. Filled markers represent the athermal system ($\tilde{D}_{t}=0$), and unfilled markers represent the thermal system ($\tilde{D}_{t}=1.0$). Data are fitted to $a_{\rho}\,\Lambda^{\nu}$ with best-fit exponents $\nu=2.00\pm 0.01$ (athermal) and $\nu=2.11\pm 0.06$ (thermal). Inset: density-dependent prefactor $a_{\rho}$ decreases with density as $(\rho-\rho_{0})^{-1}$, with $\rho_{0}\sigma^{2}=0.947\pm 0.008$ (athermal) and $\rho_{0}\sigma^{2}=0.971\pm 0.004$ (thermal).
@@
III.2 Scaling verification via numerical simulations
For this purpose, we analyze the global non-affine parameter $X=N^{-1}\sum_{i}^{N}\chi_{i}$. Its evolution reaches a steady state for $\Lambda<10$, but diverges for $\Lambda>10$ due to irreversible reorganization of particle positions. Examples of such trajectories of $X(t)$ at $\Lambda=4$ and $\Lambda=10$ are shown in Fig. 2(a). The steady-state regime is used to test the scaling hypothesis presented above.
We now consider the deviation of the global and steady-state non-affine parameter $\Delta X(\Lambda,\rho,T)=X(\Lambda,\rho,T)-X_{eq}(\rho,T)$ from equilibrium. Simulations with $N=4096$ ABPs, $\Lambda\in[0.5,8.0]$, $\rho\in[1.0,1.4]$, and $\tilde{D}_{t}=1.0$ yield a reasonable data collapse when data from different densities are plotted as $\langle\Delta X\rangle/a_{\rho}$ versus $\Lambda$. The excess non-affinity shows a scaling form $\langle\Delta X\rangle/a_{\rho}\sim\Lambda^{\nu}$ with $\nu=2.112\pm 0.006$; see Fig. 2(b). The inset in Fig. 2(b) shows the density dependence $a_{\rho}\sim(\rho-\rho_{0})^{-1}$ with $\rho_{0}\sigma^{2}=0.971\pm 0.004$, surpassing the equilibrium melting point $\rho^{eq}_{m}\sigma^{2}=0.92$ [76], reflecting active softening.
At $\Lambda\lesssim 1$, thermal fluctuations dominate over activity, making the $\Lambda$ dependence on the non-affine parameter difficult to discern. To address this, we set $\tilde{D}_{t}=0$ and explore the range $\Lambda\in[10^{-2},8]$. The results (filled symbols in Fig. 2) collapse onto a single curve across densities, resembling the $\tilde{D}_{t}=1.0$ case. Fitting the power law $a_{\rho}\,\Lambda^{\nu}$ gives $\nu=2.00\pm 0.01$, validating our predicted scaling for both thermal and athermal systems and confirming $\Delta X\sim\Lambda^{2}$ across three decades. The density-dependent prefactor $a_{\rho}\sim(\rho-\rho_{0})^{-1}$, with $\rho_{0}\sigma^{2}=0.947\pm 0.008$ (inset of Fig. 2(b) ). Though density scaling remains unchanged, $\rho_{0}\sigma^{2}$ is suppressed in the absence of thermal noise, as expected. The above analysis confirms both scaling forms presented in Eq. (8). The robustness of the scaling properties is further verified simulating larger system sizes in Appendix-B.
III.3 Probability distribution of $X$
In Figure 4, we explore the probability distributions of the global non-affine parameter $X$ for an ABP solid at $\tilde{D_{t}}=1$ and $\rho\sigma^{2}=1.1$. Figure 4a compares $P(X)$ for $N=4096$ particles across different $\Lambda$ values, showing that as $\Lambda$ increases, the mode of the distribution shifts to larger $X$ (consistent with the $\langle X\rangle\sim\Lambda^{2}$ scaling) and broadens with higher activity. This behavior capturing softening even in the absence of defect formation (as we show later, the onset of defect formation appears at $\Lambda\approx 10$). As shown in Fig. 4b for $\Lambda=4$ and $N$ ranging from 1024 to 16384, the probability distribution $P(X)$ is nearly Gaussian with variance $\sigma^{2}\sim 1/N$, consistent with the central limit theorem for the mean of uncorrelated entities. Similar scaling behavior for fluctuations in $X$ was observed before in equilibrium solids [39].
IV Analyzing structural indicators
Thus far, we have characterized the softening of a high-density solid through the non-affine parameter $X$ and its scaling with the Péclet number and density. In this section, we extend our analysis to include elastic and structural indicators, such as the shear modulus, solid order, and hexatic order, to further explore the active softening of solid. We perform this analysis using $\rho\sigma^{2}=1.1$ and $\tilde{D}_{t}=1.0$.
@@ IV.1 Shear modulus
Within linear response, the shear modulus is given by $\mathcal{G}=\sigma^{tot}_{xy}/{\cal E}_{xy}$. We compute this from numerical simulations by applying a shear strain ${\cal E}_{xy}$ and measuring the change in total shear stress $\sigma_{xy}^{\mathrm{tot}}$ accounting for trajectories displaying linear response; see Appendix-C. The stress tensor has contribution from interaction calculated using the usual virial term and swim stress due to activity. The standard virial expression due to interaction gives,
@@
$\sigma^{sw}_{\alpha\beta}=\frac{\gamma v_{0}}{D_{r}V}\sum_{i=1}^{N}\big{\langle}\dot{r}_{i,\alpha}\,u_{i,\beta}\big{\rangle}.$ (10)
The total shear stress is then calculated as $\sigma^{tot}_{xy}=\sigma^{vir}_{xy}+\sigma^{sw}_{xy}$. Appendix-C outlines details of the shear modulus calculation.
Figure 3(a) shows the change in mean shear modulus, $\cal G$, with standard errors, as a function of activity $\Lambda$. We observe a steady decrease in it from the equilibrium value indicating softening of the solid with activity. As activity increases, fraction of trajectories following linear response decreases to vanish near $\Lambda=10$; see Appendix-C. The decrease in $G$ with activity can be described by ${\cal G}={\cal G}_{0}-{\cal G}_{1}\Lambda^{2}$ with ${\cal G}_{0}=109.84$ and ${\cal G}_{1}=0.094$. To further explore this softening and the fluidization transition, we compute structural indicators, including solid order $\psi_{\mathbf{G}}$, hexatic order $\psi_{6}$, and defect fraction $\delta_{d}$.
Figure 5: The structure factor $S(\mathbf{q})$ at $\mathrm{\Lambda}=4$ (a), $10$ (b), $10.2$ (c), $10.5$ (d), $15$ (e), and $20$ (f).
@@ IV.2 Solid order
where $\rho_{\bf q}=\frac{1}{N}\sum_{i=1}^{N}\exp(i{\bf q}\cdot{\bf r}_{i})$. Plots of $S({\bf q})$ at representative $\Lambda$ values are shown in Fig.5. In a perfect triangular lattice, quasi-Bragg peaks emerge at ${\bf q}_{p}=(0,\pm 2\pi/a_{y}),(\pm 2\pi/a,\pm\pi/a_{y})$, as shown in Fig. 5(a),(b). The solid order parameter $\langle\psi_{\bf G}\rangle$ averages $S({\bf q})$ over the six quasi-Bragg peaks at ${\bf G}:=\{{\bf q}_{p}\}$. Fig. 5(b) shows the variation of $\langle\psi_{\bf G}\rangle$ with $\Lambda$. The order parameter $\langle\psi_{\bf G}\rangle$ vanishes as the solid melts. At the melting point $\Lambda=10.0$ the variance of the order parameter $\langle(\Delta\psi_{\bf G})^{2}\rangle$ peaks (inset, Fig. 3(b) ). This analysis shows that the quadratic scaling of non-affinity with activity is observed across the entire activity range before solid melting.
@@
IV.3 Hexatic order
This section examines the change in the hexatic order parameter to capture hexatic melting. Each particle in the system can be associated with a hexatic bond orientational order. For the $k$-th particle, it is defined as $\psi_{6}^{k}=(1/n)\sum_{i=1}^{n}e^{i6\theta_{kj}}$, where $\theta_{kj}$ is the bond angle between particle $k$ and its topological neighbor $j$ relative to the $x$-axis, and $n$ is the number of such neighbors [81]. In Fig. 3(c), we plot the mean amplitude of the hexatic order parameter of the system
as a function of $\Lambda$. The hexatic order diminishes with $\Lambda$, but remains significant beyond $\Lambda=10$, suggesting the system stays in the hexatic phase past the solid melting point. At the hexatic melting point $\Lambda=15$, the hexatic order variance $\langle(\Delta\psi_{6})^{2}\rangle$ peaks sharply (inset of Fig. 3(c) ).
The local amplitudes $|\psi_{6}^{k}|^{2}$ and arguments arg($\psi_{6}^{k}$) of hexatic order for typical configurations across the transition are shown in Fig. 6 as the active solid softens and melts with increasing $\Lambda$. The drops in local hexatic amplitude, linked to hexatic orientation shifts, signal topological defect formation, as detailed in the following section.

### v2

@@ III.1 Scaling analysis
Figure 2: (a) Representative time series of the system-averaged non-affinity $X(t)$ at $\Lambda=8.6$ (athermal, top) and $\Lambda=4.0$ (thermal, bottom); other parameters are given in the legends. (b) Scaled excess non-affinity $\langle X^{(a)}\rangle/\Lambda^{2}$ as a function of the active persistence $\tilde{D}_{r}^{-1}$. The data are well fit by $\langle X^{(a)}\rangle/\Lambda^{2}=a/(D_{r}+c)$, with $a=9.21\times 10^{-3}\pm 4.57\times 10^{-4}$ and $c=397.96\pm 20.57$. At large $\tilde{D}_{r}$ (low persistence), $\langle X^{(a)}\rangle$ grows approximately linearly with the persistence time $\tilde{D}_{r}^{-1}$ and saturates beyond $\tilde{D}_{r}^{-1}>c^{-1}$. (c) $\langle X^{(a)}\rangle/a_{\rho}$ versus $\Lambda$ for $N=4096$ particles at fixed $\tilde{D}_{r}=3.0$ and different densities $\rho$, where $X^{(a)}=X(\Lambda)-X_{\mathrm{eq}}$. Filled (open) symbols correspond to athermal ($\tilde{D}_{t}=0$) and thermal ($\tilde{D}_{t}=1.0$) systems, respectively. Fits to $a_{\rho}\Lambda^{\nu}$ yield $\nu=2.00\pm 0.01$ (athermal) and $\nu=2.11\pm 0.06$ (thermal). Inset: density dependence of $a_{\rho}\sim(\rho-\rho_{0})^{-1}$, with $\rho_{0}\sigma^{2}=0.947\pm 0.008$ (athermal) and $0.971\pm 0.004$ (thermal).
III.2 Scaling verification via numerical simulations
For this purpose, we analyze the system-averaged non-affinity $X=N^{-1}\sum_{i}^{N}\chi_{i}$. At $\rho\sigma^{2}=1.1$ and $\tilde{D}_{r}=3$, the time evolution $X(t)$ reaches a steady state for $\Lambda\lesssim 8$, whereas for $\Lambda\gtrsim 8$ it may diverge due to irreversible reorganization of particle positions. As we show later, the melting of the solid associated with the proliferation of topological defects occurs at a larger active speed, $\Lambda=10$; indicating that the divergence of non-affinity precedes solid melting, in agreement with earlier observations in Ref. [40].
Representative trajectories $X(t)$ at $\Lambda=4$ ($\tilde{D}_{t}=1$) and $\Lambda=8.6$ ($\tilde{D}_{t}=0$) are shown in Fig. 2(a). At high active speeds, the trajectories of $X$ exhibit large excursions that can subsequently relax back toward baseline values, indicating the healing of nascent plastic deformations. In contrast, as illustrated in the top panel of Fig. 2(a), some trajectories undergo runaway growth, signaling irreversible plastic flow. In the following, we therefore restrict our analysis to the steady-state regime $\Lambda\lesssim 8$ to test the scaling hypothesis proposed above.
We now consider the excess non-affinity induced by activity, $X^{(a)}(\Lambda,\rho,T)=X(\Lambda,\tilde{D}_{r},\rho,T)-X_{eq}(\rho,T)$. This quantity can be obtained in two equivalent ways. One approach is to simulate an active system coupled to a thermal bath ($\tilde{D}_{t}=1$) and compare it with the corresponding equilibrium system at $\Lambda=0$, thereby determining $X(\Lambda,\tilde{D}_{r},\rho,T)$ and $X_{\mathrm{eq}}(\rho,T)$ separately. Alternatively, one may perform athermal simulations with $\tilde{D}_{t}=0$, which directly isolate the contribution of activity to the system’s non-affinity.
[…]
In Fig. 2(c), we present simulations of $N=4096$ ABPs for $\tilde{D}_{r}=3.0$, $\Lambda\in[0.5,8.0]$, densities $\rho\in[1.0,1.4]$, and $\tilde{D}_{t}=1.0$, from which we extract the excess steady-state non-affinity due to activity $\langle X^{(a)}\rangle$. These results are compared with athermal simulations performed at $\tilde{D}_{t}=0$. A reasonable data collapse is obtained when results at different densities are plotted as $\langle X^{(a)}\rangle/a_{\rho}$ versus $\Lambda$. The excess non-affinity follows a scaling form $\langle X^{(a)}\rangle/a_{\rho}\sim\Lambda^{\nu}$ with exponent $\nu\approx 2.0$, as shown in Fig. 2(c). The inset of Fig. 2(c) displays the density dependence of the prefactor, $a_{\rho}\sim(\rho-\rho_{0})^{-1}$, with $\rho_{0}\sigma^{2}=0.971\pm 0.004$. This fitted threshold exceeds the equilibrium melting density $\rho_{m}^{\mathrm{eq}}\sigma^{2}=0.92$ [34], indicating activity-induced softening of the solid.
[…]
III.3 Coexistence
In Fig.3, we show the probability distributions $P(\chi)$ of local non-affinity $\chi_{i}=\chi$ for an athermal system ($\tilde{D}_{t}=0$) on log-log axes for varying active speed (panel (a) ) and persistence time (panel (b) ). In equilibrium, the distribution is described by a chi-squared form [41]. As the activity parameters increase, the slightly skewed distributions broaden, get more skewed, and develop a heavy tail, followed by the appearance of a secondary maximum indicating coexisting small and large non-affinities. Such coexistence of pronounced non-affine zones with an otherwise normal solid precedes the divergence of non-affinity, which emerges as a precursor to defect formation and eventual melting.
Figure 3: Probability distributions $P(\chi)$ of local non-affinity $\chi_{i}=\chi$ on log-log axes. (a) At fixed $\tilde{D}_{r}=3.0$, increasing active speed $\Lambda$ shifts and broadens the distributions toward larger $\chi$, with a secondary high-$\chi$ peak emerging at large $\Lambda$. (b) At fixed $\Lambda=6.0$, decreasing $\tilde{D}_{r}$ drives a crossover from unimodal to bimodal distributions, reflecting the onset of rare, large non-affine rearrangements.
In panel (a), at fixed rotational diffusion $\tilde{D}_{r}=3.0$, increasing the self-propulsion speed $\Lambda$ systematically shifts the skewed distribution toward larger values of $\chi$ and broadens it, indicating enhanced non-affine rearrangements consistent with the increase of the mean $\langle X\rangle\sim\Lambda^{2}$. At sufficiently high $\Lambda$, the distribution becomes more skewed, first developing a heavy tail at large $\chi$ and then a secondary peak, signaling the coexistence of dominant small non-affine fluctuations with rare large rearrangements.
Fig. 3(b) shows the effect of varying $\tilde{D}_{r}$ at fixed active speed $\Lambda=6.0$. For small persistence ($\tilde{D}_{r}>0.25$), the distributions remain unimodal; the low-$\chi$ tail is progressively suppressed, shifting probability weight toward larger $\chi$ as persistence increases ($\tilde{D}_{r}$ decreases). This behavior is consistent with the linear increase of $\langle X\rangle$ in the small-persistence regime. At $\tilde{D}_{r}=0.25$, the distribution undergoes a qualitative change, marked by a clear rightward shift and broadening of the peak. For $\tilde{D}_{r}\leq 0.25$, the main maximum – carrying the bulk of the probability – saturates and becomes insensitive to further increases in persistence. A long tail at large $\chi$ develops for $\tilde{D}_{r}\leq 0.25$, eventually giving rise to a secondary peak at high $\chi$, rendering the distribution bimodal. This reflects the coexistence of dominant small non-affine fluctuations with rare but substantial non-affine rearrangements.
Further, the change in the character of the probability distributions $P(\chi)$ across a broad range of $\Lambda$ and $\tilde{D}_{r}$, highlighting the emergence of coexistence, is shown in Appendix D. This detailed analysis reveals that increasing active speed broadens and shifts $P(\chi)$ toward larger values, while intermediate persistence promotes bimodality, reflecting the coexistence of small and rare large non-affine rearrangements, with very high or very low persistence restoring unimodal, tail-suppressed distributions.
III.4 Correlations
To further characterize the local properties of non-affinity, we analyze the spatial correlation
$C_{\chi}(r)=\langle\chi_{i}\chi_{j}\,\delta(r_{ij}-r)\rangle$ (8)
at steady state, where $r=|{\mathbf{r}}|$, $\chi_{i}$ denotes the local non-affinity for particle $i$, and $r_{ij}=|{\mathbf{r}}_{i}-{\mathbf{r}}_{j}|$ is the distance between particles $i$ and $j$. The normalized two-point correlation function $C\chi(|\mathbf{r}|)/C\chi(0)$ at fixed activity $\Lambda=5$ is shown in Fig. 4(a). These correlations clearly extend over longer distances as persistence increases (i.e., as $\tilde{D}_{r}^{-1}$ grows). We extract a characteristic correlation length $\xi$ using the integral definition $\xi=\int dr\,C_{\chi}(r)/C_{\chi}(0)$, which captures the correct trends, although the precise numerical value can vary depending on the specific choice of correlation length measure. Mapping $\xi$ across a range of $\Lambda$ and $\tilde{D}_{r}$ (Fig. 4(b)) reveals three distinct regimes. For high persistence, $\tilde{D}_{r}\lesssim 0.2$, $\xi$ saturates to a $\Lambda$-dependent plateau, reflecting extended correlations in persistent-active systems. In the intermediate regime, $0.2\lesssim\tilde{D}_{r}\lesssim 110$, $\xi$ decreases monotonically with $\tilde{D}_{r}$, indicating progressively shorter-ranged correlations as persistence weakens. Finally, in the low-persistence regime, $\tilde{D}_{r}\gtrsim 110$, the short correlation lengths become largely independent of activity. Increasing $\Lambda$ systematically enhances $\xi$ in the persistent-active and intermediate regimes, showing that stronger self-propulsion promotes correlations of non-affine rearrangements over longer distances, whereas in the low-persistence limit the correlation length collapses to a small value, independent of $\Lambda$.
Figure 4: Normalized spatial correlations of the local non-affinity $\chi$ (a) at $\Lambda=5$ for varying $\tilde{D}_{r}$ and (b) corresponding correlation length $\xi$ across active speeds $\Lambda$ and rotational diffusivities $\tilde{D}_{r}$. $\xi$ increases with active speed and persistence in the persistent-active and intermediate regimes, while in the low-persistence limit, correlations are short-ranged and largely $\Lambda$-independent.
Figure 5: Normalized time autocorrelation $C_{X}(\tau)$ at $\Lambda=5$ for several rotational diffusion constants $\tilde{D}_{r}$. Dashed lines indicate the reference exponential decay $\exp(-2\tilde{D}_{r}\tau)$.
In addition, we compute the two-time autocorrelation of the system-averaged non-affinity $X$, using the normalized definition
$C_{X}(\tau)=\langle\delta X^{(a)}(t)\,\delta X^{(a)}(t+\tau)\rangle/\langle[\delta X^{(a)}(t)]^{2}\rangle$ (9)
where $\delta X^{(a)}(t)=X^{(a)}(t)-\langle X^{(a)}(t)\rangle$ and the averaging is performed over time $t$. Our simulations show that $C_{X}(\tau)$ decays as $C_{X}(\tau)\sim\exp(-2\tilde{D}_{r}\tau)$, provided the system has reached a steady state in terms of non-affine fluctuations. As an example, in Fig. 5 we show numerical estimates of $C_{X}(\tau)$ at $\Lambda=5$ confirming this exponential decay for all tested values of $\tilde{D}_{r}$. This demonstrates that the dynamics of non-affinity under activity are primarily governed by the persistence of the active motion.
IV Analyzing structural indicators
Thus far, we have characterized the softening of a high-density solid through the non-affine parameter $X$ and its scaling with activity characterized by persistence and active speed, and density. In this section, we extend our analysis to include elastic and structural indicators, such as the shear modulus, solid order, and hexatic order, to further explore the active softening of solid. We perform this analysis using $\rho\sigma^{2}=1.1$, $\tilde{D}_{r}=3.0$ and $\tilde{D}_{t}=1.0$.
@@ IV.1 Shear modulus
Within linear response, the shear modulus is given by $\mathcal{G}=\sigma^{tot}_{xy}/{\cal E}_{xy}$. We compute this from numerical simulations by applying a shear strain ${\cal E}_{xy}$ and measuring the change in total shear stress $\sigma_{xy}^{\mathrm{tot}}$ accounting for trajectories displaying linear response; see Appendix-E. The stress tensor has contribution from interaction calculated using the usual virial term and swim stress due to activity. The standard virial expression due to interaction gives,
@@
$\sigma^{sw}_{\alpha\beta}=\frac{\gamma v_{0}}{D_{r}V}\sum_{i=1}^{N}\big\langle\dot{r}_{i,\alpha}\,n_{i,\beta}\big\rangle.$ (11)
The total shear stress is then calculated as $\sigma^{tot}_{xy}=\sigma^{vir}_{xy}+\sigma^{sw}_{xy}$. Appendix-E outlines details of the shear modulus calculation.
Figure 6(a) shows the change in mean shear modulus, $\cal G$, with standard errors, as a function of activity $\Lambda$. We observe a steady decrease in it from the equilibrium value indicating softening of the solid with activity. As activity increases, fraction of trajectories following linear response decreases to vanish near $\Lambda=10$; see Appendix-E. The decrease in $G$ with activity can be described by ${\cal G}={\cal G}_{0}-{\cal G}_{1}\Lambda^{2}$ with ${\cal G}_{0}=109.84$ and ${\cal G}_{1}=0.094$. To further explore this softening and the fluidization transition, we compute structural indicators, including solid order $\psi_{\mathbf{G}}$, hexatic order $\psi_{6}$, and defect fraction $\delta_{d}$.
Figure 7: The structure factor $S(\mathbf{q})$ at $\mathrm{\Lambda}=4$ (a), $10$ (b), $10.2$ (c), $10.5$ (d), $15$ (e), and 2020 (f).
@@ IV.2 Solid order
where $\rho_{\bf q}=\frac{1}{N}\sum_{i=1}^{N}\exp(i{\bf q}\cdot{\mathbf{r}}_{i})$. Plots of $S({\bf q})$ at representative $\Lambda$ values are shown in Fig.7. In a perfect triangular lattice, quasi-Bragg peaks emerge at ${\bf q}_{p}=(0,\pm 2\pi/a_{y}),(\pm 2\pi/a,\pm\pi/a_{y})$, as shown in Fig. 7(a),(b). The solid order parameter $\langle\psi_{\bf G}\rangle$ is calculated by taking an average over $S({\bf q})$ values calculated at the six quasi-Bragg peaks located at ${\bf G}:=\{{\bf q}_{p}\}$:
$\langle\psi_{\bf G}\rangle=\frac{1}{6}\sum_{p=1}^{6}\left\langle\left|\frac{1}{N}\sum_{i=1}^{N}\exp(i{\bf q}_{p}\cdot{\mathbf{r}}_{i})\right|^{2}\right\rangle$ (13)
Fig. 6(b) shows the variation of $\langle\psi_{\bf G}\rangle$ with $\Lambda$. The order parameter $\langle\psi_{\bf G}\rangle$ vanishes as the solid melts. At the melting point $\Lambda=10.0$ the variance of the order parameter $\langle(\Delta\psi_{\bf G})^{2}\rangle$ peaks (inset of Fig. 6(b) ). This analysis shows that the quadratic scaling of non-affinity with activity is observed across the entire activity range before solid melting.
@@
IV.3 Hexatic order
This section examines the change in the hexatic order parameter to distinguish hexatic melting from solid melting. Each particle in the system can be associated with a hexatic bond orientational order. For the $k$-th particle, it is defined as $\psi_{6}^{k}=(1/n)\sum_{i=1}^{n}e^{i6\theta_{kj}}$, where $\theta_{kj}$ is the bond angle between particle $k$ and its topological (Voronoi) neighbor $j$ relative to the $x$-axis, and $n$ is the number of such neighbors. In Fig. 6(c), we plot the mean squared amplitude of the hexatic order parameter of the system
as a function of $\Lambda$. The hexatic order diminishes with $\Lambda$, but remains significant beyond $\Lambda=10$, suggesting the system stays in the hexatic phase past the solid melting point, with the order vanishing near $\Lambda=20$. The hexatic melting point $\Lambda=15$ is identified by the peak in the variance of hexatic order $\langle(\Delta\psi_{6})^{2}\rangle$ (inset of Fig. 6(c) ).
@@ IV.3 Hexatic order
The local squared amplitudes $|\psi_{6}^{k}|^{2}$ and arguments arg($\psi_{6}^{k}$) of hexatic order for typical configurations across the transition are shown in Fig. 8 as the active solid softens and melts with increasing $\Lambda$. The drops in local hexatic amplitude, linked to hexatic orientation shifts, signal topological defect formation, as detailed in the following section.